# iLRM: An Iterative Large 3D Reconstruction Model

### v1

@@ 4.4 Quantitative Results
pixelSplat charatan2024pixelsplat 125 25.89 0.858 0.142 131,072 0.101
MVSplat chen2025mvsplat 12 26.39 0.869 0.128 131,072 0.047
GS-LRM* zhang2025gs-lrm 300 28.10 0.892 0.114 131,072 -
DepthSplat xu2024depthsplat 354 27.47 0.889 0.114 131,072 0.065
Gen-Den nam2024generative 28 27.08 0.879 0.120 347,072 0.224
Ours $(2,F,F)$ 171 28.17 0.894 0.115 131,072 0.025
Ours $(4,H,F)$ 185 30.01 0.918 0.098 65,536 0.027
Ours-MC $(4,H,F)$ 185 29.75 0.915 0.101 65,536 0.027
Ours $(8,H,F)$ 185 31.01 0.930 0.088 131,072 0.028
Ours-MC $(8,H,F)$ 185 30.69 0.927 0.090 131,072 0.030
Table 1: Quantitative comparisons on the RE10K zhou2018re10k dataset with various viewpoint configurations. Inference time is the reconstruction time (not rendering speed), using a single NVIDIA RTX 4090 GPU. * indicates methods for which the official code and the pretrained models are not available.
Method Views PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ # of Gaussians Time (S) Memory (GB)
MVSplat chen2025mvsplat 6 22.93 0.775 0.193 688,128 0.279 5.87
DepthSplat xu2024depthsplat 24.19 0.823 0.147 688,128 0.102 3.55
Ours $(6,H,F)$ 25.07 0.812 0.189 172,032 0.031 1.40
$(11,H,F)$ 26.47 0.852 0.155 315,392 0.051 1.59
$(24,H,F)$ 27.97 0.886 0.128 688,128 0.123 2.01
Table 2: Quantitative comparisons on the DL3DV ling2024dl3dv dataset with various view configurations. Inference time is the reconstruction time (a single NVIDIA RTX 4090), and memory consumption is measured during the reconstruction (a single forward pass).
Method ACID liu2021acid DL3DV ling2024dl3dv
@@
MVSplat chen2025mvsplat 28.15 0.841 0.147 22.65 0.737 0.191
DepthSplat xu2024depthsplat 28.37 0.847 0.141 24.28 0.813 0.147
Gen-Den nam2024generative 28.61 0.847 0.141 22.92 0.750 0.188
Ours $(2,F,F)$ 28.73 0.847 0.153 25.02 0.815 0.154
Ours $(4,H,F)$ 29.80 0.871 0.140 27.60 0.870 0.127
Ours-MC $(4,H,F)$ 29.60 0.867 0.144 27.38 0.865 0.132
Ours $(8,H,F)$ 30.51 0.886 0.131 28.98 0.897 0.112
Ours-MC $(8,H,F)$ 30.24 0.883 0.134 28.78 0.895 0.114
Table 3: Cross-dataset generalization results on the ACID liu2021acid and DL3DV ling2024dl3dv datasets. All models are trained on the RE10K zhou2018re10k training set and evaluated on each dataset without fine-tuning.
4.5 High-resolution Experiment
We present high-resolution experimental results (512×960) in Table 4 and Figure 7. Our method is compared with DepthSplat xu2024depthsplat using a 100-frame interval and 12 input images/viewpoints. The relatively small gap in inference time is that DepthSplat had to use a smaller model, as the computational cost for processing high-resolution images was significantly higher.
Method Views PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ # of Gaussians Time (S) Memory (GB)
DepthSplat xu2024depthsplat 12 21.38 0.739 0.265 5,898,240 0.504 24.93
Ours $(12,H,F)$ 23.92 0.764 0.274 1,474,560 0.415 3.52
Table 4: Quantitative comparisons on the DL3DV ling2024dl3dv dataset on high-resolution setting. Inference time and memory consumption are measured only during the Gaussian generation stage, excluding the rendering process on a single H100 GPU.
Figure 7: Qualitative comparison of novel view synthesis on the high-resolution (512$\times$960) DL3DV dataset.
@@ 4.6 Computational Costs of Training
Baseline 30.15 0.920 0.097 1.51 62.5 3.83
w/ Half Cross-attn 30.02 0.919 0.098 1.13 47.4 1.71
w/ Quater Cross-attn 29.85 0.916 0.101 0.94 39.0 0.81
Table 5: Quantitative comparisons of our mini-batch cross-attention on the RE10K zhou2018re10k dataset.
4.7 Ablations and Analysis
@@
12 layers (base) 185M 28.96 0.902 0.111
9 layers 139M 28.74 0.898 0.114
6 layers 94M 28.42 0.893 0.119
3 layers 48M 27.83 0.882 0.130
@@
Baseline 28.96 0.902 0.111
w/o token uplifting 28.69 0.898 0.116
w/o self-attention 23.29 0.753 0.222
w/ group-attention 28.76 0.899 0.115
@@
In Tab. 9, we present the ablation results on key architectural components to validate the effectiveness of each design choice. All experiments are conducted under the same configuration and settings as the model size ablation with 24 layers baseline.
1) Token uplifting. Removing the token uplifting mechanism leads to a drop in performance across all metrics compared to baseline. This validates the importance of expanding low-resolution view tokens before cross-attention with high-resolution image tokens. Without this step, the model struggles to capture fine-grained spatial correspondences, resulting a degraded reconstruction quality.
2) Self-attention. To ensure a fair comparison, we replaced all self-attention layers with cross-attention layers rather than simply removing them, maintaining a comparable parameter count. The performance dropped significantly, highlighting the essential role of self-attention in capturing global dependencies and enhancing multi-view awareness among viewpoint embeddings. Without self-attention, the model struggles to integrate contextual information across different viewpoints, resulting in poor convergence and reconstruction quality.
3) Group-attention This variant replaces the per-viewpoint cross-attention mechanism with a group-attention approach, where all viewpoint tokens and image tokens are concatenated and jointly processed through a cross-attention block (Fig. 2-(c)). Unlike our default design, group-attention introduces global interactions across all views. While this mechanism can increase the expressive capacity between multiple viewpoints, it incurs quadratic complexity with respect to the number of views. However, the increased computational cost does not yield performance gains, suggesting that separating the roles—using cross-attention for localized image-view interactions and self-attention for global refinement across viewpoints—leads to a more efficient and effective architecture.
@@ 5 Limitations
One limitation of this work is the computational bottleneck associated with self-attention across multiple input views. Although we significantly reduce the computational cost by utilizing compact viewpoint embeddings, challenges may arise as the number of input views increases considerably. In this study, aiming for scalable feed-forward 3D models, we present the first implementation of the proposed framework, a feed-forward 3D model that iteratively refines 3D representations by leveraging high-resolution image information at every layer. Further development of more scalable alternatives to self-attention, such as hierarchical and sparse attention mechanisms or other efficient processing strategies, would be a valuable direction for future research.
Additionally, our model requires known camera poses, typically obtained from structure-from-motion (SfM) Schonberger_2016_CVPR_colmap ; pan2024global methods such as COLMAP or from precisely synthesized datasets. However, these poses can be error-prone and are often difficult to acquire at scale. To enhance the practicality of our model and enable the use of giant-scale raw video datasets chen2024panda , extending it to operate in a pose-free setting ye2024no ; hong2024pf3plat ; kang2024selfsplat remains an important and open direction for future work.
@@ 6 Conclusion
Appendix A Additional Implementation Details
@@
We initialize model weights using a zero-mean normal distribution with a standard deviation of 0.02. Bias terms are omitted in all Linear and normalization layers. The model is trained using the AdamW loshchilov2017adamw optimizer with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.095$. A weight decay of 0.05 is applied to all parameters except the weights of LayerNorm ba2016layernorm . We use a cosine learning rate schedule with a peak learning rate of 2e-4 and a warmup of 2500 iterations. Our training setup largely follows the configuration proposed in zhang2025gs-lrm ; jin2024lvsm .
@@
The 8-view half-resolution viewpoint setting (8, H, F) is trained on 8 H100 GPUs with a total batch size of 256 for 50,000 iterations. Similarly, the 4-view half-resolution viewpoint setting (4, H, F) is trained on 8 RTX 4090 GPUs with a total batch size of 128 for 100,000 iterations. The mini-batch cross-attention variants were also trained with the equivalent computational budgets for each viewpoint setting using the RealEstate 10K zhou2018re10k dataset.
For training on the DL3DV dataset ling2024dl3dv , we initialize from the pretrained (8, H, F) model trained on the RE10K dataset zhou2018re10k , and finetuned on 8 H100 GPUs with a total batch size of 96 for 100,000 itertions. During training, the number of input viewpoints is randomly sampled between 6 and 11 to expose the model to varying numbers of viewpoints/images xu2024depthsplat . An additional 50,000 iterations are performed for high-resolution finetuning with varying numbers of viewpoints/images.
@@
Gaussian representations. After the final self-attention layer, the viewpoint features are decoded into Gaussian parameters using a single linear layer with an output dimension of 16. The Gaussian positions, denoted as $\mu$, consist of 5 channels: 2 for the spatial xy offset and 3 for depth, z. The final depth is obtained by averaging the 3 depth channels. Opacity ($\alpha$) is represented by a single channel. Covariance ($\Sigma$) is derived from 3 channels of scale and 4 channels of rotation. Finally, color ($c$) is represented using 3 channels. Higher-order spherical harmonics coefficients are not used in our method. The post-activation functions for each parameter follow the design of GS-LRM zhang2025gs-lrm , except for the spatial xy offset, for which we constrain the range to lie within a single pixel of viewpoint resolution. We utilize gsplat ye2025gsplat , an open-source library for Gaussian Splatting kerbl20233dgs for a rasterizer.
@@
Camera pose normalization. We normalize camera poses to align the scene into a consistent coordinate system and scale. First, we compute the average position and viewing directions (forward, down, and right) from the input camera extrinsics. These are used to build a new reference pose, which centers and aligns the scene. All camera extrinsics are then transformed into this reference frame. Finally, we scale the entire scene so that the largest camera distance is 1, ensuring the scene fits within a normalized space zhang2025gs-lrm .
@@
Appendix B Additional Architectural Details.
@@
Tokenization and normalization. After tokenizing the viewpoints and multi-view images using linear layers, both types of tokens are passed through a LayerNorm ba2016layernorm . In each cross-attention layer, only the viewpoint tokens are further processed with a pre-normalization layer. Additionally, after the query and key linear projections, both tokens are passed through an extra normalization layer, referred to as the QK-Norm henry2020qknorm .
Appendix C Additional Evaluation Details.
When we utilize more input viewpoints (more than two in RealEstate10K zhou2018re10k experiment compared to the baselines), we sample additional viewpoints/images evenly between the two endpoint indices, ensuring that these samples do not overlap with the target indices. For the DL3DV evaluation setting, we compare against the baselines using a 6-view configuration. In our 11-view configuration, we sample one additional viewpoint between every pair of consecutive baseline viewpoints. For 24 viewpoints setting, we finetuned the model with a fixed number of views for an additional 100,000 iterations.
DL3DV evaluation. For cross-dataset generalization on the DL3DV dataset, we use a baseline of 12 frames. For in-domain low-resolution evaluation, the baseline includes 50 frames in total.
Appendix D Computational Costs of Training

### v2

@@ 3.4 Results
pixelSplat (Charatan et al., 2024) 125 25.89 0.858 0.142 131,072 0.101
MVSplat (Chen et al., 2025) 12 26.39 0.869 0.128 131,072 0.047
GS-LRM* (Zhang et al., 2025) 300 28.10 0.892 0.114 131,072 -
DepthSplat (Xu et al., 2025) 354 27.47 0.889 0.114 131,072 0.065
Gen-Den Nam et al. (2025) 28 27.08 0.879 0.120 347,072 0.224
Ours $(2,F,F)$ 171 28.65 0.900 0.110 131,072 0.025
Ours $(4,H,F)$ 185 30.37 0.923 0.095 65,536 0.027
Ours-MC $(4,H,F)$ 185 30.10 0.919 0.098 65,536 0.027
Ours $(8,H,F)$ 185 31.57 0.935 0.082 131,072 0.028
Ours-MC $(8,H,F)$ 185 30.96 0.931 0.088 131,072 0.030
Table 1: Quantitative comparisons on the RE10K dataset with various viewpoint configurations. Inference time is the reconstruction time (not rendering speed), using a single NVIDIA RTX 4090 GPU. * indicates methods for which the official code and the pretrained models are not available.
Method ACID (Liu et al., 2021) DL3DV (Ling et al., 2024)
@@
MVSplat (Chen et al., 2025) 28.15 0.841 0.147 22.65 0.737 0.191
DepthSplat (Xu et al., 2025) 28.37 0.847 0.141 24.28 0.813 0.147
Gen-Den (Nam et al., 2025) 28.61 0.847 0.141 22.92 0.750 0.188
Ours $(2,F,F)$ 29.24 0.856 0.143 25.35 0.826 0.144
Ours $(4,H,F)$ 30.10 0.877 0.138 27.90 0.877 0.122
Ours-MC $(4,H,F)$ 29.90 0.873 0.141 27.68 0.881 0.127
Ours $(8,H,F)$ 30.96 0.894 0.122 29.56 0.907 0.101
Ours-MC $(8,H,F)$ 30.47 0.887 0.132 29.03 0.900 0.110
Table 2: Cross-dataset generalization results on the ACID and DL3DV datasets ($\times256$). All models are trained on the RE10K training set and evaluated on each dataset.
Method (50-frame baseline) Views PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ # of Gaussians Time (S) Memory (GB)
MVSplat (Chen et al., 2025) 6 22.93 0.775 0.193 688,128 0.279 5.87
DepthSplat (Xu et al., 2025) 6 24.19 0.823 0.147 688,128 0.102 3.55
11 24.28 0.833 0.141 1,261,568 0.170 6.01
24 22.37 0.781 0.195 2,752,512 0.371 12.39
Ours $(6,H,F)$ 25.60 0.830 0.168 172,032 0.031 1.40
$(11,H,F)$ 26.99 0.865 0.140 315,392 0.051 1.59
$(24,H,F)$ 27.38 0.882 0.126 688,128 0.123 2.01
[…]
Method (100-frame baseline) Views PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ # of Gaussians Time (S) Memory (GB)
DepthSplat (Xu et al., 2025) 12 21.38 0.739 0.265 5,898,240 - OOM
Ours $(12,H,F)$ 24.35 0.781 0.256 1,474,560 0.415 3.53
[…]
Method (full baseline; average 350-frame) views Time $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
3D-GS (30k) (Kerbl et al., 2023) 16 13min (A100) 21.20 0.708 0.264
Mip-splatting (30k) (Yu et al., 2024) 16 13min (A100) 20.88 0.712 0.274
LongLRM (Ziwen et al., 2025) 16 0.28sec (H100) 22.66 0.740 0.292
Ours $(16,H,F)$ 0.19sec (H100) 22.91 0.766 0.295
3D-GS (30k) 32 13min (A100) 23.60 0.779 0.213
Mip-splatting (30k) 32 13min (A100) 23.32 0.784 0.217
LongLRM 32 0.61sec (H100) 24.10 0.783 0.254
Ours $(32,H,F)$ 0.53sec (H100) 24.30 0.803 0.257
LongLRM10 32 37sec (A100) 25.60 0.826 0.233
Ours10 $(32,H,F)$ 7sec (A100) 25.67 0.844 0.230
Ours50 $(32,H,F)$ 34sec (A100) 26.46 0.862 0.210
[…]
Figure 5: Qualitative comparison of novel view synthesis on the RE10K dataset ($\times256$).
[…]
@@ 3.5 Computational Costs of Training
Baseline 30.39 0.923 0.095 1.51 62.5 3.83
w/ Half Cross-attn 30.25 0.922 0.096 1.13 47.4 1.71
w/ Quater Cross-attn 30.08 0.919 0.098 0.94 39.0 0.81
Table 6: Quantitative comparisons of our mini-batch cross-attention on the RE10K dataset.
@@
12 layers (base) 185M 29.24 0.907 0.109
9 layers 139M 29.01 0.903 0.112
6 layers 94M 28.68 0.898 0.116
3 layers 48M 28.04 0.887 0.126
@@
Baseline 29.24 0.907 0.109
w/o iter. refinement 28.58 0.893 0.127
w/o resolution decoupling 28.47 0.891 0.123
w/o token uplifting 28.90 0.901 0.113
@@ 3.6 Ablations and Analysis
In Tab. 9, we present the ablation results on key architectural components to validate the effectiveness of each design choice. All experiments are conducted under the same configuration as the model size ablation with 12 layers baseline. Further ablation studies are provided in the supplementary material.
1) Iterative refinement. The cross-attention blocks in our model keep providing visual (image) into the viewpoint tokens as part of the iterative refinement process. To validate our design choice, we conduct an ablation study in which only the first layer uses cross-attention, while the remaining layers use self-attention. The baseline consists of 12 layers, each composed of a cross-attention followed by a self-attention block, whereas the variant uses a single cross-attention layer followed by 23 self-attention layers. The table below shows that our consecutive cross-attention with image features plays a critical role in refining the viewpoint embeddings especially comparing LPIPS metric.
2) Resolution decoupling. We perform an ablation study where low-resolution images (matched to the viewpoint resolution) are used as visual inputs for the cross-attention layers. Compared to the baseline, which uses high-resolution images, this variant shows a performance drop, confirming the effectiveness of our decoupling strategy. It enables the use of high-resolution features while maintaining a compact 3D representation.
3) Token uplifting. Removing the token uplifting mechanism leads to a drop in performance across all metrics compared to baseline. This validates the importance of expanding low-resolution view tokens before cross-attention with high-resolution image tokens. Without this step, the model struggles to capture fine-grained spatial correspondences, resulting a degraded reconstruction quality.
@@ 4 Limitations
One limitation of this work is the computational bottleneck associated with self-attention across multiple input views. Although we significantly reduce the computational cost by utilizing compact viewpoint embeddings, challenges may arise as the number of input views increases considerably. In this study, aiming for scalable feed-forward 3D models, we present the first implementation of the proposed framework, a feed-forward 3D model that iteratively refines 3D representations by leveraging high-resolution image information at every layer. Further development of more scalable alternatives to self-attention would be a valuable direction for future research.
Additionally, our model requires known camera poses, typically obtained from structure-from-motion (SfM) (Schonberger & Frahm, 2016; Pan et al., 2024) methods such as COLMAP or from precisely synthesized datasets. However, these poses can be error-prone and are often difficult to acquire at scale. To enhance the practicality of our model and enable the use of giant-scale raw video datasets (Chen et al., 2024b), extending it to operate in a pose-free setting (Ye et al., 2024; Hong et al., 2024; Kang et al., 2025) remains an important and open direction for future work.
@@ 5 Conclusion
6 Acknowledgements
This research was partially supported by a grant of the Korea-US Collaborative Research Fund (KUCRF), funded by the Ministry of Science and ICT and Ministry of Health & Welfare, Republic of Korea (grant number: RS-2024-00468417), the Institute of Information and communications Technology Planning & Evaluation (IITP) grant funded by the Korean government(MSIT) (No. RS-2024-00457882, AI Research Hub Project, RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University), and RS-2025-02653113, High-Performance Research AI Computing Infrastructure Support at the 2 PFLOPS Scale).
Appendix A Related Works
A.1 Feed-forward 3D Gaussian Splatting
Feed-forward 3D Gaussian Splatting (Charatan et al., 2024; Chen et al., 2025; Xu et al., 2025; Nam et al., 2025; Zhang et al., 2025; Szymanowicz et al., 2024b; Tang et al., 2025; Xu et al., 2024b) capitalizes on robust priors learned from extensive datasets to estimate Gaussian primitive parameters and synthesize novel view images using sparse input data. PixelSplat (Charatan et al., 2024) and LatentSplat (Wewer et al., 2024) predict Gaussians from image features using an epipolar line sampling method to enhance geometric accuracy, while MVSplat (Chen et al., 2025) and MVSGaussian (Liu et al., 2024) construct cost volumes through a plane-sweep stereo approach. In a further development, Flash3D (Szymanowicz et al., 2024a) and DepthSplat (Xu et al., 2025) introduce a pre-trained depth estimation model (Piccinelli et al., 2024; Yang et al., 2024), which improves the robustness of the spatial positions of 3D Gaussians. In contrast, GS-LRM (Zhang et al., 2025) and Long-LRM (Ziwen et al., 2025) minimize reliance on explicit 3D priors by leveraging large-scale data-driven priors.
Iterative refinements. Our work is also closely related to recent methods that adopt iterative refinement strategies, such as G3R (Chen et al., 2024e) and Gen-Den (Nam et al., 2025). Both utilize actual gradients to update their representations more precisely. While promising, these approaches require additional computational burden for rendering multiple images per training iteration, and relying solely on gradients may risk overlooking valuable information present in the raw input images. Nonetheless, exploring how to incorporate gradient information remains an interesting direction for future work.
A.2 3D Representations from Embeddings
Inspired by previous generative approaches (Goodfellow et al., 2014; Karras et al., 2019; Chan et al., 2022), recent works (Hong et al., 2023; Chen et al., 2024a; Flynn et al., 2024) have investigated the synthesis of 3D representations directly from learnable embeddings, guided by input image supervision. This paradigm leverages the expressive capacity of latent spaces to encode rich geometric priors, which act as structural templates that guide the reconstruction process. Such approaches offer notable flexibility, allowing rendering from arbitrary viewpoints and adaptation to varying space scales and camera poses. However, both LRM (Hong et al., 2023) and Lara (Chen et al., 2024a) are limited to object-centric representations, restricting scalability to complex scenes involving multiple objects or large spatial layouts. The recently proposed Quark (Flynn et al., 2024) also utilizes learnable embeddings to fuse visual cues from multiple images, demonstrating compelling results, but its representation is confined to the target view (Xu et al., 2024a; Liu et al., 2024; Jin et al., 2024), lacking an explicit and persistent 3D reconstruction.
Appendix B Additional Implementation Details
@@
We initialize model weights using a zero-mean normal distribution with a standard deviation of 0.02. Bias terms are omitted in all Linear and normalization layers. The model is trained using the AdamW (Loshchilov & Hutter, 2017) optimizer with hyperparameters $\beta_{1}=0.9$ and $\beta_{2}=0.095$. A weight decay of 0.05 is applied to all parameters except the weights of LayerNorm (Ba et al., 2016). We use a cosine learning rate schedule with a peak learning rate of 2e-4 and a warmup of 2500 iterations. Our training setup largely follows the configuration proposed in (Zhang et al., 2025; Jin et al., 2024).
@@
For the RealEstate10K (RE10K) (Zhou et al., 2018) dataset, the 8-view half-resolution viewpoint setting (8, H, F) is trained on 8 H100 GPUs with a total batch size of 256 for 50,000 iterations. Similarly, the 4-view half-resolution viewpoint setting (4, H, F) is trained on 8 RTX 4090 GPUs with a total batch size of 128 for 100,000 iterations. The mini-batch cross-attention variants were also trained with the equivalent computational budgets for each viewpoint setting. Lastly, the 2-view full-resolution viewpoint setting (2, F, F), which serves as the reference point, is trained on 8 H100 GPUs for 200,000 iterations.
[…]
@@
Gaussian representations. After the final self-attention layer, the viewpoint features are decoded into Gaussian parameters using a single linear layer with an output dimension of 16. The Gaussian positions, denoted as $\mu$, consist of 5 channels: 2 for the spatial xy offset and 3 for depth, z. The final depth is obtained by averaging the 3 depth channels. Opacity ($\alpha$) is represented by a single channel. Covariance ($\Sigma$) is derived from 3 channels of scale and 4 channels of rotation. Finally, color ($c$) is represented using 3 channels. Higher-order spherical harmonics coefficients are not used in our method. The post-activation functions for each parameter follow the design of GS-LRM (Zhang et al., 2025), except for the spatial xy offset, for which we constrain the range to lie within a single pixel of viewpoint resolution. We utilize gsplat (Ye et al., 2025), an open-source library for Gaussian Splatting Kerbl et al. (2023) for a rasterizer.
@@
Camera pose normalization. We normalize camera poses to align the scene into a consistent coordinate system and scale. First, we compute the average position and viewing directions (forward, down, and right) from the input camera extrinsics. These are used to build a new reference pose, which centers and aligns the scene. All camera extrinsics are then transformed into this reference frame. Finally, we scale the entire scene so that the largest camera distance is 1, ensuring the scene fits within a normalized space (Zhang et al., 2025).
@@
Appendix C Additional Architectural Details
@@
Tokenization and normalization. After tokenizing the viewpoints and multi-view images using linear layers, both types of tokens are passed through a LayerNorm (Ba et al., 2016). In each cross-attention layer, only the viewpoint tokens are further processed with a pre-normalization layer. Additionally, after the query and key linear projections, both tokens are passed through an extra normalization layer, referred to as the QK-Norm (Henry et al., 2020).
Appendix D Additional Evaluation Details
When we utilize more input viewpoints (more than two in RealEstate10K (Zhou et al., 2018) experiment compared to the baselines), we sample additional viewpoints/images evenly between the two endpoint indices, ensuring that these samples do not overlap with the target indices. For cross-dataset generalization on the DL3DV dataset, we use a baseline of 12 frames.
In wide-baseline setting, every 8th image in the sequence is reserved for the test split, while K-means clustering on camera positions and viewing directions is applied to the remaining images to select input views that ensure wide scene coverage.
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
Baseline 29.24 0.907 0.109
w/o self-attention 23.33 0.755 0.220
w/ group-attention 29.02 0.904 0.112
w/ random init. 28.90 0.902 0.112
w/ LR-feature init. 28.35 0.894 0.121
Figure 11: Additional ablations.
Appendix E Addition Ablations on Model Architecture
We provide additional ablation studies and analyses in Tab. 11 under the same configuration as the ablations on model architecture in main script. All variants are trained under half-resolution 4 viewpoints setting $(4,H,F)$, with a batch size of 16 on a single RTX 4090 GPU.
1) Self-attention. To ensure a fair comparison, we replaced all self-attention layers with cross-attention layers rather than simply removing them, maintaining a comparable parameter count. The performance dropped significantly, highlighting the essential role of self-attention in capturing global dependencies and enhancing multi-view awareness among viewpoint embeddings. Without self-attention, the model struggles to integrate contextual information across different viewpoints, resulting in poor convergence and reconstruction quality.
2) Group-attention This variant replaces the per-viewpoint cross-attention mechanism with a group-attention approach, where all viewpoint tokens and image tokens are concatenated and jointly processed through a cross-attention block. Unlike our default design, group-attention introduces global interactions across all views. While this mechanism can increase the expressive capacity between multiple viewpoints, it incurs quadratic complexity with respect to the number of views. However, the increased computational cost does not yield performance gains, suggesting that separating the roles—using cross-attention for localized image-view interactions and self-attention for global refinement across viewpoints—leads to a more efficient and effective architecture, which is also validated as Alternating-Attention in VGGT (Wang et al., 2025a).
3) Different initialization. We also investigate the different initialization methods of scene representation. For the random initizliation, we used a learnable embedding initialized with zero mean and 0.02 standard deviation, whereas for the LR feature variant, we used features extracted from low-resolution images. In the PSNR training curve, the LR feature variant rises more quickly in the early stages but is later surpassed by the random initialization variant. We believe this is because, in our iterative cross-attention architecture, high-resolution image features and camera information are continually provided by the cross-attention blocks. As a result, the learnable embedding can offer more flexible and informative parameters for guiding iterative updates, whereas the LR image features may introduce redundant and less discriminative information that limits long-term performance gains. Moreover, the use of LR features may bias the early attention stages toward coarse processing, which can hinder the model’s ability to fully refine fine details in later stages.
Appendix F Computational Costs of Training